### v1

Abstract
Large language models (LLMs) have revolutionized natural language processing. However, effectively incorporating complex and potentially noisy user interaction data remains a challenge. To address this, we propose User-LLM, a novel framework that leverages user embeddings to contextualize LLMs. These embeddings, distilled from diverse user interactions using self-supervised pretraining, capture latent user preferences and their evolution over time. We integrate these user embeddings with LLMs through cross-attention and soft-prompting, enabling LLMs to dynamically adapt to user context. Our comprehensive experiments on MovieLens, Amazon Review, and Google Local Review datasets demonstrate significant performance gains across various tasks. Notably, our approach outperforms text-prompt-based contextualization on long sequence tasks and tasks that require deep user understanding while being computationally efficient. We further incorporate Perceiver layers to streamline the integration between user encoders and LLMs, reducing computational demands.
1 Introduction
Large language models (LLMs) have revolutionized the field of natural language processing (NLP) Brown et al. (2020); Chowdhery et al. (2023); OpenAI (2023); Touvron et al. (2023); Anil et al. (2023); Google (2023). With their ability to learn and adapt from massive amounts of textual data, LLMs offer significant opportunities for user modeling and personalization. By analyzing user interactions and understanding user preferences, LLMs can be leveraged to power recommendations Liu et al. (2023b); Lyu et al. (2023); Ji et al. (2023), language generation, summarization Liu et al. (2023d); Basyal & Sanghvi (2023), and question answering Kojima et al. (2022); Wei et al. (2023) in ways that are highly relevant and engaging to users. This paves the way for creating more personalized and context-aware language-based applications and services.
Figure 1: Illustration of two different mechanisms to incorporate user interaction history data to LLMs. (1) User history data are expressed in natural language and fed to LLMs via text prompt; (2) User embeddings, distilled from user history data, are used to contextualize LLMs.
User interactions represent a rich source of behavioral data generated from a user’s engagement with digital systems. Spanning a wide range from textual input, search queries, media consumption (e.g., videos watched or rated), to social media activities, navigation patterns, location visits, and more, these interactions hold valuable insights for user modeling. One straightforward approach to leveraging this data with LLMs is to directly finetune LLMs on textual components, using a user’s interaction history as the text prompt. However, user interaction data is often complex, spanning multiple journeys with sparse data points, various interaction types (multimodal), and potential noise or inconsistencies. This complexity can hinder an LLM’s ability to identify and focus on the most relevant patterns. Moreover, effective personalization often requires a deep understanding of the context and latent intent behind user actions, which can pose difficulties for LLMs trained predominantly on vast, surface-level language corpora. Additionally, user interaction data, such as extended histories, can be very lengthy. Processing and modeling such long sequences (e.g., a year’s worth of history) with LLMs can strain computational resources, making it practically infeasible. Addressing these challenges is key to unlocking the full potential of LLMs in user modeling and personalization.
To address the inherent complexities and limitations of leveraging raw user interaction data with LLMs, we propose a novel approach centered around user embeddings, as illustrated in Fig. 1. These compressed representations, distilled from diverse and noisy user interactions, effectively capture the essence of a user’s behavioral patterns and preferences across various interaction modalities. By contextualizing the LLM with user embeddings during finetuning or inference, we aim to
  • Enhance its ability to identify relevant patterns navigating complexity and noise,
  • Facilitate understanding and adaptation to the latent intent, dynamic context, and temporal evolution behind user actions,
  • Mitigate the computational demands of processing extensive interaction histories by working with condensed representations.
This approach empowers LLMs with a deeper understanding of users’ historical patterns and latent intent, enabling LLMs to tailor responses and generate personalized outcomes.
Our approach consists of two key phases: generating high-quality user embeddings and contextualizing LLMs with these user embeddings. In phase one, we pretrain a Transformer-based encoder on user interaction data, utilizing self-supervised learning to capture behavioral patterns across multiple interaction modalities. We use a multi-feature autoregressive Transformer to generate embeddings that capture long-range dependencies and contextual relationships within sequential data while handling multimodal user data effectively. In phase two, we integrate user embeddings with an LLM during finetuning using cross-attention, where the LLM’s intermediate text representations attend to the output embeddings from the pretrained user encoder, enabling dynamic context injection (similar to Flamingo Alayrac et al. (2022)).
We present a comprehensive study across three public datasets and various tasks. Our findings conclusively demonstrate that the embedding-based approach successfully mitigates the limitations associated with directly finetuning LLMs on raw textual user interaction data. The approach shows great potential to enhance performance in several application domains, including user understanding, personalized recommendations, and text generation.
Our contributions are four-fold:
  • We introduce User-LLM, a versatile framework that leverages user embeddings distilled from diverse modalities (e.g., video watch history, ratings, location visits) to contextualize LLMs. It dynamically incorporates user preferences and behaviors from various interaction modalities, enhancing LLM understanding and personalization capabilities while supporting various encoder architectures and multimodal fusion mechanisms.
  • We rigorously assess User-LLM against state-of-the-art baselines, demonstrating significant performance improvements, particularly in tasks requiring deep user understanding and long context inputs. Our approach excels in personalization while maintaining computational efficiency, outperforming traditional text-prompt-based contextualization.
  • User-LLM offers flexible training strategies. Notably, finetuning only the user encoder while keeping the LLM frozen effectively contextualizes the LLM and outperforms LoRA-based text-prompt LLM tuning (which also preserves original LLM weights). This highlights User-LLM’s adaptability, effectively contextualizing the LLM without altering its core, maintaining language modeling and generalization abilities.
  • We provide in-depth analyses of cross-attention and soft-prompt approaches for integrating user embeddings, exploring the potential of Perceiver layers for further efficiency gain. Extensive ablation studies offer practical guidance on the impact of different User-LLM components and design choices.
2 Related Work
2.1 User Modeling
Dual encoders and self-supervised learning are widely employed for user models and recommendation systems, providing transferable representations as user understanding and personalization signals Covington et al. (2016b; a); Sountsov & Sarawagi (2016); Volkovs et al. (2017); Chidambaram et al. (2018); Gillick et al. (2018); Yanga et al. (2018); Ma et al. (2018); Yi et al. (2019); Gillick et al. (2019); Yang et al. (2020); Jiang et al. (2020); Ning et al. (2021). Other self-supervised learning approaches such as contrastive learning Yao et al. (2021); Xie et al. (2022); Chen et al. (2022b); Xia et al. (2023) and graph representation learning Cai et al. (2023); Yang et al. (2023) have proven to learn high-quality representations from massive but noisy user interaction data. Recently, pre-trained BERT model Devlin et al. (2018) has been applied to sequential recommendation systems. For instance, Bert4Rec Sun et al. (2019) and U-Bert Qiu et al. (2021) leverage pre-training and fine-tuning approaches to model users’ behavior in a self-supervised way. Our work explores the potential of integrating user representations with the emerging capabilities of LLMs to further improve user modeling and personalization.
2.2 Language Model based Personalization
The emergence of LLMs has propelled generative models to demonstrate superior natural language generation capabilities compared to discriminative models. There has been extensive work on different ways to format user historical interactions as prompts and to leverage LLMs to generate recommendations across various tasks Petrov & Macdonald (2023); Kang et al. (2023); Xu et al. (2023b); Liu et al. (2023b); Lyu et al. (2023); Ji et al. (2023); Li et al. (2023); Wu et al. (2023a). However, directly incorporating a user’s entire history into an LLM’s context window is often impractical or overly expensive. In contrast to the aforementioned approaches, we propose to leverage user embeddings to contextualize LLMs.
A recent study Doddapaneni et al. (2024) explored LLM personalization using user embeddings derived from user activity text. In contrast, we utilize user activity tokens for a more efficient approach. In addition, we employ sophisticated fusion techniques and conduct a comprehensive empirical evaluation across multiple datasets and tasks to provide a thorough assessment of our method’s capabilities.
2.3 Long Context in LLMs
Long context is crucial for models to incorporate long-term user data and behavior patterns. There has been extensive work to enable long context for LLMs. One line of work is to extend context window by remapping the positional encodings of the long context to the original window size Chen et al. (2023a); Jin et al. (2024), or by exposing relevant additional contexts available to a subset of attention layers Tworkowski et al. (2023). Another direction is to distill the information in the long context to fit in the existing window by using a separate retrieval model Xu et al. (2023a) or a memory bank Wang et al. (2023) etc., or to compress the information by training special tokens Ge et al. (2023); Mu et al. (2023); Chevalier et al. (2023). Other approaches include modified attention computation Ding et al. (2023); Liu et al. (2023a); Chen et al. (2023b) and linear attention frameworks such as structured state space models Gu et al. (2023); Gu & Dao (2023); Tay et al. (2022). Our method tackles long context by using a single token to represent each user event. The extracted representations align with the LLM text embedding space. Therefore, our model is compatible with existing LLM-based techniques and applications.
2.4 Multimodal LLM
Early multimodal LLMs such as CLIP Radford et al. (2021) and ImageBind Girdhar et al. (2023) relied on aligning image and text representations. Subsequently, fusion techniques leveraging cross-attention (e.g., Flamingo Alayrac et al. (2022), Coca Yu et al. (2022)) and soft prompts (e.g., PaLI Chen et al. (2022a), Palm-E Driess et al. (2023)) emerged. Recent works, such as NextGPT Wu et al. (2023b), OneLLM Han et al. (2023), and Anymal Moon et al. (2023), explored unified frameworks for diverse input modalities. Additionally, end-to-end training of multimodal models has gained attention, as seen in Gato Reed et al. (2022) and Gemini Gemini Team Google (2023). Inspired by these advances, User-LLM focuses on efficient LLM contextualization with user understanding capabilities.
3 User-LLM
Figure 2: Overview of User-LLM. Left: Multimodal Autoregressive Transformer encoder pretraining. Right: LLM contextualization with user embeddings. Features from a user’s interaction history are encoded by the Autoregressive User Encoder and then integrated into the language model via cross-attention.
This section introduces the User-LLM framework, a two-stage approach for contextualizing LLM with user embeddings, as illustrated in Fig.2. First, a pretrained user encoder generates user embeddings. Second, these embeddings are integrated into the LLM through cross-attention or soft-prompt techniques, providing the LLM additional context and guidance to generate personalized responses.
3.1 Embedding Generation
User-LLM leverages a Transformer-based encoder to generate user embeddings from ID-based feature sequences. Consider a scenario where user data contains two modalities, $a$ and $b$. Each modality has its own vocabulary. Items from each modality are mapped to integer IDs and have distinct embedding representations. To create a combined representation, aligned items $a_{i}$ and $b_{i}$ are fused into a single embedding $f_{i}$. The sequence of fused embeddings serves as input to the user encoder for generating user embeddings.
Autoregressive Encoder In this work, we employ an Autoregressive Transformer as the user encoder. As illustrated in Fig. 2 (left), this model takes a sequence of user activities, where each activity is described by multiple modalities (e.g., item name, rating, category). Individual features are embedded separately, then concatenated and processed by a Transformer decoder. The output is projected back to the original feature spaces, where modality-specific softmax layers calculate logits. Note that input embeddings and output softmax layers share weights.
The autoregressive design enables the model to predict upcoming tokens based on previous ones, using cross-entropy loss for optimization. Given a sequence of $N$ input tokens, the Autoregressive Transformer generates $N$ corresponding embeddings. These embeddings serve as user context for LLMs, enabling personalized response generation.
Our framework is adaptable, supporting various user encoders. We’ve experimented with alternatives like Dual Encoders and achieved competitive results.
3.2 LLM Contextualization with User Embedding
3.2.1 Model Architecture
User-LLM integrates user embeddings with LLMs using cross-attention (see Fig. 2). The output embeddings from the pretrained user encoder are cross-attended with intermediate text representations within the LLM, similar to how Flamingo Alayrac et al. (2022) works. Note that our framework remains adaptable to different integration mechanisms including prepending embeddings as soft-prompt Lester et al. (2021).
3.2.2 Efficiency
Compared to text-prompt-based methods, User-LLM offers substantial efficiency gains. First, it leverages pretrained weights of both the user encoder and LLM, improving model convergence speed with less trainable parameters. Additionally, User-LLM condenses user activities into dense representations, requiring only a single token per event. This frees up the LLM’s context window, leading to improved inference efficiency. Furthermore, User-LLM leverages cross-attention, maintaining user embeddings at a smaller dimension than the LLM’s model dimension and eliminating the need for projection operations. This design choice significantly reduces computational overhead during both training and inference.
To further optimize inference efficiency, User-LLM integrates Perceiver Jaegle et al. (2021) units into its projection layers. Perceiver is a Transformer-based architecture that utilizes a trainable latent query to extract relevant information from input representations through cross-attention. In User-LLM, Perceiver effectively compresses the user embeddings into a compact format using a latent query with shorter sequence length. This compression reduces the number of tokens needed to represent the user history, further freeing up the LLM’s context window. This, along with Perceiver’s ability to distill insights from noisy contexts, makes User-LLM ideal for practical applications with extensive user histories.
3.2.3 Training Strategies
User-LLM offers flexible training strategies to tailor its performance for diverse use cases. To thoroughly explore User-LLM’s capabilities, we primarily investigated four strategies targeting different model components:
  • Full: Finetuning the entire model (LLM, user encoder, and projection layers) enables maximum parameter adaptation to user interactions, revealing the upper bounds of User-LLM’s personalization potential.
  • Enc: Finetuning the user encoder and projection layers (LLM frozen) offers an efficient LLM contextualization approach while preventing the LLM from overfitting to specific user data. This maintains the LLM’s general language capabilities.
  • LoRA: Finetuning the LLM with LoRA (Low-Rank Adaptation) Hu et al. (2022), along with the user encoder and projection layers, offers parameter efficiency preventing catastrophic forgetting and maintaining the LLM’s core knowledge.
  • Proj: Finetuning only the projection layers (LLM and user encoder frozen) assesses the minimal parameter tuning required for effective LLM contextualization.
4 Experiments
4.1 Datasets and Tasks
Dataset Train Examples Test Example Seq Len Item
MovieLens20M 13,821,405 81,970 50 Movie
Google Review 3,387,375 80,080 50 Place
Amazon Review 357,258 5,756 50 Product
Table 1: Dataset train / test split sizes, sequence lengths, and item features.
We evaluated our approach on three widely recognized datasets. Table 1 summarizes their statistics after processing. Notably, Amazon Review dataset is well known for its high sparsity and variability and contains a significantly smaller number of training and evaluation examples.
  • MovieLens20M [^1] https://grouplens.org/datasets/movielens/ This is a widely used benchmark dataset Harper & Konstan (2015) for evaluating personalization and recommendation algorithms. Interaction features are movie name, genre, and rating.
  • Google Local Review Dataset [^2] https://datarepo.eng.ucsd.edu/mcauley_group/gdrive/googlelocal/ Li et al. (2022); Yan et al. (2023) This dataset contains review information on Google Maps up to Sep 2021. We used New York’s data in our experiments. Interaction features are place name, place category, rating, and review.
  • Amazon Review Dataset [^3] https://nijianmo.github.io/amazon/index.html He & McAuley (2016) This is a series of product review datasets crawled from Amazon.com. We used the ”Movie_and_TV” category for evaluation. Interaction features are product title, product category, rating, and review summary, which is a short opinion summary written by the user.
We generated training and test examples by applying a sliding window over each user’s feature sequences (sorted by timestamps). This creates inputs containing N items, with the subsequent item as the label. To ensure model evaluation consistency, the final example (representing the most recent interaction) is used as the test example for each user and is never included in the training set.
To address the diverse evaluation needs, we accessed model performance on three distinct types of tasks:
  • Next item prediction Given a historical sequence of items, the model predicts the subsequent item. For example, predicting the next movie a user watches based on a sequence of previously viewed movie IDs.
  • Favorite genre or category prediction The model predicts a user’s favorite genre or category based on a sequence of items (e.g., predicting favorite movie genre given a sequence of movie IDs). The favorite is the one with the highest frequency in the user’s history. This simple task serves as a proof-of-concept, demonstrating User-LLM’s ability to extract and understand user preferences from historical interactions.
  • Multimodal review generation This task extends beyond simple predictions. Here, the model must generate actual reviews, utilizing multimodal input features (e.g., movie ID, genre, and rating). These features are encoded to create embeddings that guide the generation of a user’s potential review for a given item.
4.2 Baselines and Experiment Setup
To assess the effectiveness of User-LLM, we compare it against the following baselines:
  • Dual Encoder(DualEnc) An embeddings-based two-tower model that is widely used for large-scale recommendation systems and natural language tasks.
  • Bert4Rec The state-of-the-art language model powered recommendation model BERT4Rec Sun et al. (2019). To ensure fair comparison, Bert4Rec employs similar transformer parameters as our user encoder.
  • TextPrompt(TP) An LLM contextualized with user history via text-prompt-based finetuning on the raw text user history.
Our experiments primarily used an Autoregressive Transformer for embedding generation, with ablation studies comparing performance against a Dual Encoder architecture. For co-training, we explored four different training strategies described in Section 3.2.3 and evaluated the impact of using a pretrained encoder versus a randomly initialized one.
For all text-prompt-based LLM finetuning and User-LLM experiments, we kept the total number of examples being processed by the model the same for comparison. For Dual Encoder and Bert4Rec baselines, we did hyper-parameter tuning and report the best performance results.
Our experiments utilized PaLM-2 XXS Anil et al. (2023) as the pretrained LLM. The user encoder is a $6$-layer $128$-dimensional transformer with $8$ attention heads. For Perceivers, we constructed the module with $6$ cross-attention layers and a query dimension of $16$. We used $rank=8$ for all LoRA tuning experiments.
4.3 Overall Performance Comparison
This subsection compares model accuracy across various tasks and baselines. All reported User-LLM experiments utilize pretrained encoders and finetune all three components (encoder, projection layers, and LLM) during cotraining, unless stated otherwise.
4.3.1 Next Item Prediction
Table 2 shows that User-LLM outperforms the two non-LLM baselines on MovieLens and Google Local review datasets for next item prediction. However, Bert4Rec excels on the sparse Amazon Review dataset, likely due to its bidirectional architecture, which enables more robust representation learning from limited training examples. In contrast, our autoregressive (unidirectional) approach relies on larger datasets for effective user modeling. This suggests that Amazon Review dataset may not provide sufficient information for the autoregressive approach to learn effectively for next item prediction tasks.
Dataset Rec Baseline User-LLM
DualEnc Bert4Rec
MovieLens20M @1 0.044 0.038 0.054
@5 0.135 0.135 0.164
@10 0.206 0.158 0.243
Google Review @1 0.005 0.005 0.015
@5 0.012 0.019 0.052
@10 0.023 0.033 0.071
Amazon Review @1 0.021 0.034 0.037
@5 0.026 0.051 0.047
@10 0.031 0.062 0.051
Table 2: User-LLM v.s. DualEnc & Bert4Rec baselines for next item prediction. Note that we use full finetune for User-LLM.
Fig. 3 compares User-LLM and text-prompt-based LLM finetuning on MovieLens20M (more results in Appendix).
Varying Context Length When trained on sequences with predefined lengths (e.g., 50, 100, or 200 items) and all parameters tunable (solid lines in Fig. 3), User-LLM shows slightly lower performance than text-prompt-based finetuning with shorter sequences (50 items). However, User-LLM significantly outperforms the text-prompt-based approach when sequences become longer. This is likely because next-item prediction heavily relies on the LLM’s ability to memorize specific item names (e.g., movie titles). Text prompt excels in this type of task by providing the LLM with numerous names or titles to memorize. As sequences become longer, they introduce diverse and potentially noisy information, making pure memorization less effective. It is a known limitation that LLMs are less effective with long input context Liu et al. (2023c). In contrast, User-LLM’s ability to effectively distill relevant user information from these complex long sequences enhances LLM’s understanding and leads to better performance.
With increasing input sequence length, both User-LLM and text-prompt-based approach experience performance degradation. This is likely due to the significantly reduced number of training examples available (13.8M, 10.6M, and 7.0M for lengths 50, 100, and 200, respectively). While the autoregressive encoder impacts model performance due to insufficient training data, User-LLM’s performance decline is notably less severe compared to the text-prompt-based approach.
Computation Efficiency Notably, the computation cost and memory requirements of text-prompt-based approaches become increasingly prohibitive with increasing input sequence length. Table 3 highlights User-LLM’s computational efficiency. The user encoder distills ID-based user activity sequences into compact user embeddings, enabling the LLM to process a fixed-length $32$-token input query regardless of input sequence length. Conversely, the text-prompt-based approach requires LLM input token lengths to be scaled proportionally with the input sequence.
Specifically, let the number of LLM parameters be $N$, encoder parameters be $N_{e}$, and Perceiver parameters be $N_{p}$. As $N_{e}\ll N$ and $N_{p}\ll N$, we can ignore computational costs associated with the encoder and Perceiver. Assuming batch size $B$, training steps $S$, and input length (in tokens) $L$, the total number of training tokens is $D=BSL$. Applying heuristic $FLOPs\approx 6ND$ Kaplan et al. (2020), User-LLM shows up to 78.1X FLOPs reduction when the sequence length reaches 200. This demonstrates that User-LLM provides a more computationally efficient and practically feasible solution for long sequence modeling.
Figure 3: User-LLM v.s. TextPrompt(TP) baseline for next item prediction on MovieLens20M (varying sequence lengths, unfrozen or frozen LLM). Unfrozen LLM: full finetune for TP model, training strategy Full for User-LLM. Frozen LLM: LoRA parameters tuning for TP model, training strategy Enc for User-LLM.
Len50 Len100 Len200
# Tokens User-LLM 32 32 32
TP 700 1350 2500
FLOPs Reduction 21.9X 42.2X 78.1X
Table 3: LLM input token counts for User-LLM v.s. TextPrompt(TP) and FLOPs reduction achieved by User-LLM. FLOPs reduction refers to FLOPs(TP) / FLOPs(User-LLM).
Frozen LLM User-LLM excels in contextualizing LLMs while preserving their original knowledge. As described in Section 3.2.3, User-LLM allows the LLM to remain frozen (e.g., training strategy Enc), preventing catastrophic forgetting issues. Fig. 3 (dotted lines) demonstrates that User-LLM with a frozen LLM surpasses the accuracy of text-prompt-based LoRA tuning. This observation suggests that performance gains in text-prompt-based full finetuning may largely result from overfitting, leading to significant knowledge loss within the LLM.
Furthermore, User-LLM with a frozen LLM also outperforms models with unfrozen LLMs (both User-LLM and text-prompt-based approach). This highlights the value of the foundation knowledge encoded in LLMs for improving downstream task performance. By utilizing user embeddings, User-LLM enables efficient and stable LLM contextualization, effectively mitigating the drawbacks of text-prompt-based methods.
4.3.2 Deep User Understanding
Even though user encoders are pretrained on next item prediction tasks, the extracted user embeddings encapsulate generalized user context. Leveraging these embeddings, User-LLM demonstrates great generalization capability to diverse personalization tasks, as shown in Table 4. These tasks require deep user understanding, demonstrating User-LLM’s effectiveness in understanding user intent and preferences from interaction data.
User-LLM outperforms the Dual Encoder baseline, which is trained to predict genre/category ID based on a sequence of user history item IDs. Despite its encoder’s pretraining on next-item prediction, User-LLM surpasses the specialized Dual Encoder across all three datasets. Compared to the TextPrompt baseline, User-LLM excels on Amazon Review and achieves similar performance on MovieLens and Google Local Review, with a significantly smaller context window (16 times smaller), emphasizing its efficiency.
User-LLM also demonstrates its strength in the multimodal task of review generation based on the item name, category, and user rating. It exceeds the TextPrompt baseline on both Amazon Review and Google Local Review datasets. We attribute this success to User-LLM’s ability to leverage dense representations of user history, an advantage particularly evident in tasks with long inputs where text-prompt-based LLM finetuning often struggles. Note that Dual Encoder cannot handle such text generation tasks.
Dataset Task Baseline User-LLM
DualEnc TP
MovieLens 20M FavGenre 0.779 0.788 0.785
Google Review FavCat 0.812 0.887 0.862
ReviewG - 10.20 11.72
Amazon Review FavCat 0.855 0.885 0.890
ReviewG - 22.82 26.38
Table 4: Favorite genre/category generation and review generation. Metrics: Recall@1 for favorite genre/category prediction tasks and ROUGE-Lsum for review generation.
4.4 Efficiency
Parameter efficiency Embedding-LLM requires fewer tunable parameters to achieve competitive performance. Table 5 shows the model accuracy across multiple training strategies. Notably, the Enc approach (tuning encoder and projection layers and freezing the LLM) achieved comparable task accuracy to full finetuning while requiring significantly fewer tuned parameters. This is because it effectively leverages the pretrained weights of both the user encoder and LLM, minimizing the need for extensive parameter updates during training. Thus, our approach offers a parameter-efficient solution for personalizing LLMs, reducing computational demands and facilitating model tuning on resource-constrained devices.
Dataset Task Full LoRA Enc Proj
MovieLens 20M NextItem 0.054 0.053 0.052 0.040
FavGenre 0.785 0.784 0.783 0.750
Google Review NextItem 0.015 0.016 0.015 0.015
FavCat 0.862 0.844 0.844 0.615
ReviewG 11.72 11.64 11.76 9.61
Amazon Review NextItem 0.037 0.028 0.028 0.014
FavCat 0.890 0.887 0.885 0.850
ReviewG 26.38 24.56 24.30 19.04
Table 5: Comparison between different training strategies. Metric: Recall@1.
Inference efficiency Furthermore, User-LLM is inference efficient. Compared to text-prompt-based approaches that require multiple tokens to represent each event in the user history, our embedding-based approach effectively condenses event information into dense representations, requiring only a single token per event. The integration of Perceiver further reduces the total number of tokens for long user histories, resulting in even faster inference speeds and making our approach ideal for real-time applications. Table 6 shows that User-LLM and its Perceiver variant maintain competitive performance while requiring a smaller number of tokens and, as a result, significantly reduce the computation complexity in the attention modules in LLM.
Task User-LLM User-LLM +Perceiver
Full Enc Full Enc
NextItem 0.054 0.052 0.054 0.054
FavGenre 0.785 0.783 0.784 0.784
Tokens 50 16 (-68%)
Table 6: User-LLM and its Perceiver variant effectively reduce the number of embedding tokens while maintaining competitive performance. Experiments are done on MovieLens 20M. Metric: Recall@1.
4.5 Ablation Study
Dataset Recall TextPrompt User-LLM
TP10 TP50 AR50 AR50TP10
MovieLens 20M @1
0.049
0.060
0.054
0.059
@5
0.142
0.172
0.164
0.173
@10
0.211
0.251
0.243
0.252
Google Review @1 0.021 0.021
0.015
0.021
@5
0.060
0.061
0.052
0.061
@10
0.082
0.086
0.071
0.082
Amazon Review @1
0.038
0.034
0.037
0.041
@5
0.047
0.042
0.047
0.051
@10
0.050
0.047
0.051
0.055
Table 7: Experiments on short-term text prompt long-term embedding. TP10: TextPrompt baseline with $10$ recent user events. TP50: TextPrompt baseline with $50$ recent user events. AR50: With $50$ user events as encoder inputs. AR50TP10: With $50$ user events as encoder inputs and $10$ most recent events in text format as LLM inputs.
Long-term and short-term user context The design of User-LLM allows it to integrate both long-term and short-term user context. We can feed long-term user history into the user encoder to generate compact user embeddings while prompting LLM with short-term user interactions. As shown in Table 7, User-LLM generally achieves better results when combining long-term and short-term user context for next item prediction tasks. Using 50 user activities for embedding generation and 10 most recent activities for prompting LLM, User-LLM (AR50TP10) outperforms TextPrompt with only short-term context (TP10) by providing long-term user information and outperforms TextPrompt with long-term context (TP50) by extracting meaningful user representations efficiently. Compared to AR50, in which User-LLM only uses long-term user data, short-term context provides the model (AR50TP10) with additional natural language information about item names. This leads to improved next item name prediction while preserving computational efficiency.
Benefit of pretraining We study the benefits of pretraining user encoders for downstream tasks. Results presented in columns 3 and 4 in Table 8 demonstrate that models utilizing pretrained encoders consistently outperform those with randomly initialized encoders across all evaluated tasks. This finding suggests that pretraining enables the encoder to capture latent user preference effectively, providing valuable context to the LLM during the cotraining process.
Soft Prompt v.s. Cross Attention We evaluate the performance of an alternative encoder-LLM integration strategy, soft-prompt Lester et al. (2021), within User-LLM framework, in which we prepend user embeddings as soft prompt tokens to the language model and meanwhile also learn an additional task prompt. We set the task prompt’s length to $10$ and utilized a pretrained Autoregressive encoder in our experiment. As shown in the last column of Table 8, the cross-attention-based approach generally outperforms the soft-prompt approach. This advantage is particularly significant in the two review generation tasks, suggesting that the cross-attention mechanism enables LLMs to better extract and utilize the rich user information encoded within user embeddings. This highlights the value of cross-attention for tasks demanding a nuanced understanding of human intent.
Dataset Task User-LLM
cross-attention soft prompt DualEnc -Xatten
PT RD
MovieLens 20M NextItem 0.054 0.041 0.051 0.044
FavGenre 0.785 0.778 0.787 0.775
Google Review NextItem 0.015 0.015 0.013 0.014
FavCat 0.862 0.853 0.822 0.797
ReviewG 11.72 11.43 9.06 11.68
Amazon Review NextItem 0.037 0.026 0.031 0.020
FavCat 0.890 0.882 0.894 0.854
ReviewG 26.38 23.12 25.80 23.83
Table 8: Experiments on different model fusing architectures, encoders, and the benefit of pretraining. Report Recall@1 for next item prediction and favorite genre/category prediction, ROUGE-Lsum for review generation. Full finetuning is used throughout the experiments. PT: Pretrained encoder. RD: Randomly initialized encoder. soft-prompt: Connect user embeddings with LLM via soft-prompt. DualEnc-Xatten: Use a dual encoder as the user encoder and cross-attend it to LLM.
Different Encoders We experiment with an alternative Transformer-based encoder, Dual Encoder (which serves as a baseline in this work), within the User-LLM framework. This architecture features two separate user and label towers to generate embeddings. The similarity between user and label embeddings is used for loss computation via a softmax layer during pretraining. We only the use user tower when integrating with LLMs. User-LLM with Dual Encoder performed worse than with AR encoder (Table 8). One reason is that Dual Encoder only outputs a single embedding for each user, providing less context to LLM compared to the AR encoder. Additionally, while Autoregressive Transformer benefits from loss calculation across all input tokens, Dual Encoder relies on a single label per input sequence. This limitation becomes more pronounced on sparse datasets like Amazon Review, leading to a larger performance gap.
5 Conclusion and Future Work
In this paper, we introduced User-LLM, a framework for contextualizing LLMs through user embeddings. These embeddings, derived from self-supervised pretraining on diverse user interactions, capture hidden user preferences and their evolution. By integrating these embeddings with LLMs through cross-attention and soft-prompt, User-LLM empowers LLMs to adjust dynamically to user contexts.
Our comprehensive evaluation across MovieLens, Amazon Review, and Google Local Review datasets demonstrated significant performance improvements in various tasks. User-LLM showed competitive performance compared with non-LLM baselines and text-prompt-based LLM personalization techniques, particularly in handling long sequences and understanding users deeply. User-LLM’s computational efficiency and ability to preserve LLM knowledge further make it a highly suitable approach for real-world user understanding applications.
Future research could delve into optimizing user embedding generation through advanced pretraining techniques, investigating the alignment between user embeddings and the language model space for deeper user context understanding, and training User-LLM on a diverse range of tasks to enhance its generalization abilities and adaptability across a broader spectrum of user scenarios. By addressing these aspects, User-LLM has the potential to become an even more robust and versatile framework for user modeling and LLM personalization.
Appendix A Encoder Architectures
User-LLM uses a Transformer-based user encoder to generate user embeddings from ID-based feature sequences. To effectively fuse information from multimodal user interaction data, we explore and compare two fusion architectures: early fusion and late fusion. Following the example shown in Fig. 4, each item in the user interaction timeline has three modalities: name, rating, and category. Each modality (or feature) has its own vocabulary. Items from each modality are mapped to an integer ID and have their own embedding representations. In an early fusion architecture, the $i_{th}$ item’s three modalities (name, rating, and category) are combined into a fused embedding $f_{i}$. The sequence of fused embeddings is then fed into the user encoder to generate user embeddings. On the other hand, a late fusion architecture employs three separate feature encoders to process each modality independently. The resulting embeddings are combined to a single user embedding at a later stage using fusion layers.
Figure 4: Encoder architectures supported in User-LLM: early fusion and late fusion user encoders.
Autoregressive Transformer As described in Section 3.1, User-LLM employs an Autoregressive Transformer as the early fusion encoder. Illustrated in Fig. 2, this model receives a sequence of user activities, where each activity is represented by multiple modalities (name, rating, and category). Individual features are first embedded, then concatenated and processed by a Transformer decoder. The Autoregressive Transformer Encoder generates a user embedding for each input item in the user’s timeline.
Dual Encoder The late-fusion encoder we investigated is a dual encoder architecture with separate ’user’ and ’label’ towers. The ’user tower’ utilizes Transformers and processes a series of input tokens, while the ’label tower’ processes the subsequent token (label) within the sequence. Each encoder generates an embedding representation of its respective input, and the similarity between these embeddings is calculated for loss computation via a softmax layer. When integrating with LLMs, we utilize only the ’user tower’ for user embedding generation. With the default setting, a Dual Encoder generates a single user embedding from an input sequence. The number of generated embeddings can be adjusted through the configuration of the fusion layers.
Autoregressive Transformers and Dual Encoders offer different advantages for user embedding generation. Autoregressive Transformers excel at capturing long-range dependencies and contextual relationships within sequential data. However, they produce a fixed number of output embeddings (equal to the input sequence length), potentially introducing inefficiencies in the later LLM integration stage, especially when the sequence length is long. Furthermore, training these models can be computationally intensive.
Compared to Autoregressive Transformers, Dual Encoders offer several advantages. The separate feature encoders within user towers allow independent processing of features, handling features with varying sequence lengths or misaligned timestamps and potentially enhancing information extraction. The configurable fusion layers provide flexibility in the number of generated embeddings. Additionally, Dual Encoders support diverse pretraining tasks beyond next item prediction, as user and label towers can process different features (e.g., name for the user tower and category for the label tower). However, Dual Encoders may be less effective at modeling the complex sequential patterns often present in multimodal user data. The optimal encoder choice depends on dataset characteristics, the desired level of contextual understanding, and a balance between computational cost and potential performance gain. Refer to Table 8, Table 12, and Table 13 for performance comparison between the two encoder architectures.
Appendix B Integration Approaches
We explored the integration of user embeddings and LLMs in User-LLM through two approaches: cross-attention and soft-prompt. In the cross-attention method, the output embeddings from the pretrained user encoder are cross-attended with the LLM’s intermediate text representations, similar to how Flamingo Alayrac et al. (2022) works. On the other hand, the soft-prompt approach leverages user embeddings from the pretrained encoder as a soft prompt for the LLM. This soft prompt is prepended to the regular text prompt embedding, providing additional context about the user’s interests and preferences. For a comparison of cross-attention and soft-prompt approaches, refer to Table 8 and Table 14.
Appendix C Training Hyperparameters
In Table 9, we show training hyperparameters of User-LLM, including learning rate, weight decay, batch size, and number of training steps. In addition, User-LLM used the cosine learning rate decay across all experiments and used the first $10\%$ of training steps for linear warmup.
Dataset Task Learning rate Weight decay Batch size Training steps
MovieLens 20M
EncoderPretrain
1e-1
1e-1
32,768
10,000
NextItem
1e-3
1e-3
8,192
10,000
FavGenre
1e-3
1e-3
1,024
5,000
Google Review
EncoderPretrain
1e-4
1e-1
16,384
10,000
NextItem
1e-3
1e-1
2,048
5,000
FavCat
1e-3
1e-1
2,048
5,000
ReviewG
1e-3
1e-1
1,024
5,000
Amazon Review
EncoderPretrain
1e-4
1e-1
12,800
7,000
NextItem
1e-3
1e-1
8,192
10,000
FavCat
1e-3
1e-1
1,024
5,000
ReviewG
1e-3
1e-1
2,048
5,000
Table 9: Training hyperparameters.
Appendix D More Results
D.1 Next Item Prediction
Table 11 and Table 10 show results comparing text-prompt-based approach and User-LLM in next item prediction tasks.
Dataset Recall TP User-LLM
Google Review @1 0.021 0.015
@5 0.061 0.050
@10 0.086 0.071
Amazon Review @1 0.034 0.037
@5 0.042 0.047
@10 0.047 0.051
Table 10: User-LLM v.s. TextPrompt(TP) baseline for next item prediction on Google Local Review and Amazon Review datasets. Note that we use full finetune for both User-LLM and TP.
Table 11: Varying sequence length experiments on MovieLens20M with frozen LLM or unfrozen LLM. Unfrozen LLM: full finetune for TextPrompt(TP) model, training strategy Full for User-LLM. Frozen LLM: LoRA parameters tuning for TextPrompt(TP) model, training strategy Enc for User-LLM.
Recall Unfrozen LLM Frozen LLM
Len50 Len100 Len200 Len50 Len100 Len200
TP User-LLM TP User-LLM TP User-LLM TP User-LLM TP User-LLM TP User-LLM
@1 0.060 0.054
0.034
0.037
0.018
0.020
0.023
0.053
0.017
0.040
0.011
0.031
@5 0.172 0.164
0.101
0.118
0.058
0.067
0.071
0.164
0.053
0.127
0.034
0.102
@10 0.251 0.243
0.151
0.181
0.090
0.106
0.108
0.244
0.080
0.198
0.053
0.159
Table 11: Varying sequence length experiments on MovieLens20M with frozen LLM or unfrozen LLM. Unfrozen LLM: full finetune for TextPrompt(TP) model, training strategy Full for User-LLM. Frozen LLM: LoRA parameters tuning for TextPrompt(TP) model, training strategy Enc for User-LLM.
D.2 Autoregressive Encoder v.s. Dual Encoder
Table 12 compares Autoregressive Encoder and Dual Encoder on long-term and short-term user context tasks. Refer to Columns 4 and 12 in Table 13 for the performance of these two encoders when utilizing a pretrained encoder and using Full training strategy.
Dataset Recall AREnc-Xatten DualEnc-Xatten
AR50 AR50TP10 DE50 DE50TP10
MovieLens 20M @1 0.054 0.059
0.044
0.051
@5 0.164 0.173
0.135
0.151
@10 0.243 0.252
0.206
0.221
Google Review @1
0.015
0.021
0.014
0.020
@5
0.052
0.061
0.046
0.055
@10
0.071
0.082
0.066
0.078
Amazon Review @1 0.037 0.041
0.020
0.030
@5 0.047 0.051
0.024
0.038
@10 0.051 0.055
0.026
0.040
Table 12: Experiments on short-term text prompt long-term embedding with two encoder architectures: Autoregressive Encoder and Dual Encoder. AR50: With $50$ user events as Autoregressive Encoder inputs. AR50TP10: With $50$ user events as Autoregressive Encoder inputs and $10$ most recent events in text format as LLM inputs. DE50: With $50$ user events as Dual Encoder inputs. DE50TP10: With $50$ user events as Dual Encoder inputs and $10$ most recent events in text format as LLM inputs.
D.3 Training Strategies and Encoder Pretraining
Table 13 shows the performance of Autoregressive encoder based User-LLM under different training strategies. It also compares results between using a pretrained user encoder and a randomly initialized user encoder in the encoder-LLM integration stage.
Dataset Metric AREnc-Xatten(User-LLM) DualEnc-Xatten
Full LoRA Enc Proj Full + PT
PT RD PT RD PT RD PT RD
NextItem MovieLens20M
R@1
0.054
0.041
0.053
0.043
0.052
0.041
0.040
0.002
0.044
R@5
0.164
0.133
0.164
0.136
0.164
0.133
0.134
0.008
0.142
R@10
0.243
0.203
0.244
0.205
0.243
0.204
0.203
0.013
0.218
Google Review
R@1
0.015 0.015 0.016 0.015 0.015 0.015 0.015 0.015 0.014
R@5
0.050
0.052 0.051
0.050
0.051 0.051
0.050
0.042
0.046
R@10
0.071 0.071
0.070
0.070
0.070
0.070
0.070
0.057
0.066
Amazon Review
R@1
0.037
0.023
0.028
0.018
0.028
0.022
0.014
0.000
0.020
R@5
0.047
0.031
0.035
0.026
0.034
0.029
0.019
0.000
0.024
R@10
0.051
0.037
0.038
0.029
0.036
0.030
0.021
0.0003
0.026
FavGenre (FavCat) MovieLens20M
R@1
0.785
0.778
0.784
0.779
0.783
0.778
0.750
0.484
0.775
R@5
0.996
0.994
0.996
0.994
0.996
0.994
0.993
0.931
0.994
R@10
0.999
0.999
0.999
0.999
0.999
0.999
0.999
0.991
0.999
Google Review
R@1
0.862 0.853
0.844
0.785
0.844
0.775
0.615
0.142
0.797
R@5
0.951 0.944
0.938
0.915
0.937
0.911
0.855
0.505
0.920
R@10
0.964 0.956
0.948
0.941
0.949
0.936
0.909
0.660
0.944
Amazon Review
R@1
0.890
0.882
0.887
0.884
0.885
0.886
0.850
0.223
0.854
R@5
0.934 0.948
0.917
0.931
0.917
0.930
0.918
0.663
0.912
R@10
0.944 0.958
0.919
0.934
0.919
0.932
0.923
0.793
0.934
ReviewG
Google Review
ROUGE
11.72
11.43
11.64
11.55
11.76
11.56
9.61
8.05
11.68
Amazon Review
ROUGE
26.38
23.12
24.56
25.00
24.30
25.18
19.04
17.30
23.83
Table 13: User-LLM experiments results with different training strategies and different user encoder architectures. PT: Pretrained user encoder. RD: Randomly initialized user encoder.
D.4 Cross-attention v.s. Soft-prompt
Table 14 compares the performance of two encoder-LLM integration strategies: cross-attention and soft-prompt.
Dataset Metric cross-attention (User-LLM) soft-prompt
Full LoRA Enc Proj Full LoRA Enc Proj
NextItem MovieLens20M R@1 0.054 0.053 0.052 0.040 0.051 0.036 0.0047 0.0025
Google Review R@1 0.015 0.016 0.015 0.015 0.013 0.008 0.012 0.008
Amazon Review R@1 0.037 0.028 0.028 0.014 0.031 0.010 0.003 0.003
FavGenre (FavCat) MovieLens20M R@1 0.785 0.784 0.783 0.750 0.787 0.787 0.781 0.692
Google Review R@1 0.862 0.844 0.844 0.615 0.822 0.729 0.694 0.259
Amazon Review R@1 0.890 0.887 0.885 0.850 0.894 0.877 0.831 0.356
ReviewG Google Review ROUGE 11.72 11.64 11.76 9.61 9.06 9.04 8.75 7.76
Amazon Review ROUGE 26.38 24.56 24.30 19.04 25.80 22.86 20.36 15.12
Table 14: User-LLM experiments results for cross-attention & soft-prompt with different training strategies (with pretrained Autoregressive encoders). Report Recall@1 for next item prediction and favorite genre/category prediction, ROUGE-Lsum for review generation.
Appendix E Review Generation Examples
Table 15 shows some examples from Amazon/Google review generation tasks. From these examples, we can observe personalized responses generated using User-LLM exhibit greater alignment with user preferences and backgrounds compared to unpersonalized responses. While discrepancies remain between personalized responses using User-LLM and the ground truth, User-LLM demonstrates the ability to produce reviews that reflect specific user characteristics and inclinations more effectively.
Input Prompt Ground Truth Response Personalized Response Unpersonalized Response
“Write a review for The Firm - Aerobic Body Shaping.”
“Save Your Money.”
“Terrible Waste of Time.”
“Aerobic Workout.”
“Write a review for A.I. - Artificial Intelligence.”
“A pure movie masterpiece of a film by Steven Speilberg. You will either love or hate this movie, but it will make you think.”
“If you are into science fiction and you are looking for a movie that is not only good but also educational and fun, ’A.I.”
“A.I. - Artificial Intelligence.”
“Write a review for The Family that Preys.”
“Four Stars.”
“Four Stars.”
“A great thriller!”
“Write a review for Green Street Hooligans 2.”
“Hmmmmmm…”
“Not for everyone.”
“A good sequel.”
“Help me write a review for Blue with a rating of 5.0.”
“(Translated by Google) Very delicious and the service staff are very enthusiastic (Original) 非常好吃,服务人员很热忱.”
“(Translated by Google) Good food (Original) 好吃.”
“Great food and service.”
“Help me write a review for The Home Depot with a rating of 2.0.”
“(Translated by Google) The service is terrible (Original) El servicio esta pésimo.”
“(Translated by Google) Very bad service (Original) Muy mal servicio.”
“No help.”
Table 15: Review generation examples.